OpenStudy (anonymous):

Who would like to have a crack at this problem?: (5a^3)(4a^2)

OpenStudy (anonymous):

20a^13/12

OpenStudy (anonymous):

5a x 4a = 20a 2/3 + 3/2 = 13/12 Answer : 20a^13/12

OpenStudy (marissalovescats):

20a^5. When multiplying powers, you add. a^2+a^3=a^5. 4*5 is 20. So 20a^5.

OpenStudy (marissalovescats):

Beca\[b ^{n}timesb ^{m}=b ^{n+m}\]use